a conical vessel 11cm high having a top of 50 mm diameter is filled completely to its rim with water. some spherical pellets, each of 5 mm diameter are dropped into the vessel & then 2/5 of the water overflows.how many lead pallets were dropped in the vessel?
Answer by Cromlix(4381)

Volume of Conical vessel = 1/3*Pi*r^2*h
50mm = 5.0cm
Volume = 1/3 * Pi * 2.5^2 * 11
Volume = 71.995cm^3
Volume of a sphere = 4/3*Pi*r^3
= 4/3*Pi*0.25^3
= 0.0654cm^3
2/5 of the water overflows.
2/5 * 71.995cm^3 overflows
=21.798cm^3 overflows.
so x * 0.0654 = 21.798
x = 21.798/0.0654
x = 333.3
So 333 spherical pellets were dropped in the vessel.
Hope this helps.
:-)